OpenStudy (curiousreader7):

Find the derivative: g(x)=log2sqr.root(x). I've used the log. function and will comment the form I'm at now, but what next?

OpenStudy (curiousreader7):

Find the derivative of: \[g(x)=\log_{2}\sqrt{x} \] And this is where I'm at: \[g'(x)=\frac{ \frac{ 1 }{ 2 }x^{-1/2}}{ (\ln 2)\sqrt{x} }\]

OpenStudy (curiousreader7):

@agent0smith ....Have a bit of free time?

OpenStudy (campbell_st):

why not use change of base and make it a base e log function... then differentiate

OpenStudy (campbell_st):

so using change of base you get \[\log_{2} (x^{\frac{1}{2}}) = \frac{\ln(x^{\frac{1}{2}})}{\ln(2)}\]

OpenStudy (campbell_st):

remember ln(2) is a constant

OpenStudy (curiousreader7):

Oh my gosh I would have never thought of that

OpenStudy (curiousreader7):

@campbell_st , how would you differentiate that?

OpenStudy (welshfella):

use the chain rule

OpenStudy (1davey29):

Personally, I'd use the properties of logarithms to turn \[\log_{2} (x ^{1/2})=\frac{ 1 }{ 2 }\log_{2} (x) \] Then I would differentiate it into \[\frac{ 1 }{ 2x \ln(2) }\]

OpenStudy (welshfella):

yes that's an easier way